OpenStudy (anonymous):

Factor by grouping: xy - 2y + 2x - 4

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Factor like terms: y(x - 2) + 2(x - 2) Then, group common factors of x - 2 to get: (y + 2)(x - 2)

OpenStudy (anonymous):

thnks:)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

hey can you help with Factor the GCF and then factor the resulting binomial: 50x3 - 128x

OpenStudy (anonymous):

2x(5x+8)(5x-8) here ya go

OpenStudy (anonymous):

im sorry im going to be a bug but i dont understand any of this but i need to graduate can u do this one Factor the GCF first and then factor the trinomial: 2x2 + 20x + 50

OpenStudy (anonymous):

PLEASE CAN U HELP S.C ME

OpenStudy (anonymous):

s.c?

OpenStudy (anonymous):

1st factor—2: = 2x² - 20x + 50 = 2(x² - 10x + 25) 2nd & 3rd factors: x² - 5x = - 25 + (- 5)² x² - 5x = - 25 + 25 (x - 5)² = 0 x - 5 = 0 Answer: 2(x - 5)²

OpenStudy (anonymous):

this is the last one you said ^